OpenStudy (unimatix):

Integral (y^2)/(25+y^2) dy

OpenStudy (unimatix):

I'm tempted to cancel out the y^2 but I feel like that is wrong?

OpenStudy (freckles):

\[y^2+25-25=y^2\]

OpenStudy (irishboy123):

+1 @freckles then a tan sub

OpenStudy (solomonzelman):

yes \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{y^2}{y^2+25}~dy=\int\limits_{~}^{~}\left(1-\frac{25}{y^2+25}\right)~dy}\) \(u=5\tan(y)\) \(du=5\sec^2(y)\)